a.)

 N2H4 and H2O2 react according to the equation below.

 N2H4(g)+2H2O2(g)--> N2(g)+ 4H2O(g)

Using the information given in the table, determine the accepted enthalpy of reaction for the following reaction:

Delta H*/KJmol-1  N2H4(g)= +75 H2O2(g)=-133 H2O(g)=-242



b.)

some mean bond enthalpies are given below, based on the enthalpy of reaction calculated in part a, what is the mean bond enthalpy in kj/mol for the O-O bond?

Mean bond enthalpy in KJ/mol-1 N-H=388 N-N=163 N=N=944 H-O=463


2 H--O--O--H --> N=N+4 H--O--H


1
Expert's answer
2021-02-10T01:46:24-0500

a. ∆H= Hp - Hr

∆H= [4(-242)] - [75+2(-133)]

∆H= -968+191=-777KJ/mol


b. Let O-O bond energy be x

N2H4= 1N-N + 4N-H = 163 + 4(388) = 1715KJ

2H2O2= 2[2H-O + 1O-O]= 2[2(463) + 1(x)] = 2x + 1852

N2= 1N-N = 944KJ

4H2O= 4[2H-O]= 4[2(463)]= 3704KJ


Now,

∆H= Hp - HR

-777= [944 + 3704] - [1715 + 1852 + 2x]

-777= 1081 - 2x

2x= 1081 + 777

2x= 1858

x= 1858/2 = 929KJ


Therefore,

Bond energy of O-O bond= 929KJ